# A Momentum Balance Correction to the Non-Conservative One-Fluid Formulation in Boiling Flows using Volume-of-Fluid

### v1

@@ 5.3 Two-Dimensional Film Boiling
The increase in vapor volume $V$ normalized by the initial vapor volume $V_{0}$ and the average Nusselt number, Nu, on the bottom wall, defined as in Eq. (47) [17], are shown in Figure 20. Moreover, grid convergence is assessed with six mesh sizes (i.e., M1 to M6), summarized in Table 6. Up to 0.2 s, the results agree well with the reference data and are fairly independent of mesh resolution. This corresponds to the initial growth of the vapor film before the instability accelerates and the bubble forms. For reference, Figure 21 shows the evolution of the mass-averaged vapor rise velocity and its center of mass obtained with M6 (i.e., the quantities are averaged over the vapor volume due to $\rho_{G}$ being constant). After that, substantial differences in $V/V_{0}$ and Nu are observed across different mesh resolutions and, more importantly, across different numerical frameworks.
@@
At first glance, one may claim that the model is reproducing the solution from Guo et al. [59]. However, upon closer inspection of the bubble shape, this is not the case. Instead, a shift in the time scale linked to the bubble formation, i.e., instability growth, is observed. The shape of the bubble is compared in Figure 22 against benchmark solutions at $t=0.42$ s [60, 17] and at $t=0.43$ s [59], and its growth is clearly slower. The plots also show the results with mesh M6 at $t=0.46$ s, which agree much better with the reference data. This delay between 0.03 s and 0.04 s is consistent with the shift in the $V/V_{0}$ and Nu profiles shown in Figure 20. When it comes to the numerical approach, the present work is close in methodology to Boyd and Ling [17] and considers the recent state-of-the-art in multiphase flow modeling. Some differences revolve around the conservative/non-conservative treatment of the momentum equation and the implementation of the volume dilatation due to phase change, but convergence of the numerical solution is observed with similar grid resolutions to M6. In fact, one could argue that earlier works are somewhat under-resolved [59, 60].
@@
Regardless of the exact onset, the numerical solution is very similar to previous literature. Thus, the differences must arise from the flow dynamics during the growth of the vapor film and the necking process leading to the bubble formation. In particular, these are affected by the specific numerical framework and are not related to convergence issues or to capturing the thermal layer as in Section 5.2. Figure 23 presents various snapshots of the solution at $t=0.15$ s, $t=0.25$ s, $t=0.35$ s and $t=0.45$ s obtained with M6, focusing on the contours of pressure, velocity magnitude, temperature and mass flux across the interface. It becomes immediately clear that a pressure spike occurs across the interface, which scales with $\dot{m}^{\prime}$. This is a result of a momentum imbalance introduced by the discrete viscous term whereby “numerical” regularization fails. That is, a numerical pressure jump is introduced due to the calculation of $\nabla\mathbf{u}$ in the one-fluid framework, augmented by a high fluid viscosity. The thin region of the vapor film closer to the wall shows the largest momentum imbalance, while the error is much less significant along the bubble’s edge due to vapor cooling and the decrease in $\dot{m}^{\prime}$. This regularization failure occurs independently of the type of mean used to average the viscosity, i.e., arithmetic vs. harmonic.
As depicted in Figure 23, the lateral pressure gradient that drives the vapor flow and bubble growth is a result of the varying film thickness and neck formation process. Because $\dot{m}^{\prime}$ varies along the interface, the error introduced by the imbalance also does, modifying the pressure field in the vapor locally and non-uniformly. Thus, the dynamics are easily affected, such as the spurious currents around the interface, and could explain the slower bubble growth with the current formulation. Note that despite the pressure spike across the interface scales with $\sim(\Delta x)^{-1}$, the solution seems to converge but to a wrong pressure field. This is visualized in Figure 24, showing the pressure extracted along $x=17$ mm at $t=0.45$ s.
@@
In this configuration, the impact of the momentum balance corrections from the Stefan flow shift and the addition of $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime}}$ are less pronounced than in previous sections given that the viscous imbalance dominates. However, some differences are still observed that suggest that corrections in the treatment of the viscous term to improve consistency may help bring the solution closer to recent reference data [17] and reduce uncertainty. For example, Figure 25 shows a slight change in the shape of the bubble along its edge by considering or not the momentum balance corrections. The shape corresponds to the snapshot at $t=0.45$ s obtained with M6.
@@ 6 Conclusions
A two-step correction has been proposed consisting of the addition of: (1) two body forces within the context of the CSF model [20], namely $\mathbf{f}_{NC}$ to correct or cancel the effects of an ill-defined convective term and $\mathbf{f}_{\dot{m}^{\prime}}$ to recover the exact momentum jump due to phase change [12], and (2) a shift of the Stefan flow after advecting the interface to obtain an intermediate velocity field before solving the momentum equation. The latter effectively addresses the momentum jump induced by the interface displacement in the non-conservative formulation [14]. Various tests have shown that these modifications successfully impose the correct momentum balance in viscous flows with low viscosity, recovering analytical solutions and reducing oscillations in the pressure field as the interface moves across grid cells. Moreover, the improvements are also evident in dynamic flows where the momentum balance corrections are seemingly negligible and have direct implications on the dynamic behavior of the flow. For example, bubble rise velocities may differ (see Section 5.2). However, for highly viscous flows, the discretization of the viscous term can induce sharp pressure spikes across the interface and affect the momentum balance and dynamics of the flow, regardless of whether the conservative or non-conservative form of the momentum equation is used (see Section 5.3).
@@
It is in these scenarios that inconsistencies among different flow solvers are highlighted. The introduction of phase change is not trivial and requires a thorough examination of the impacts of a particular interface capturing scheme or discretization of the governing equations. Otherwise, different solvers will produce solutions similar in nature but with critical differences that may affect the particular analysis, e.g., the shift in time scales seen in film boiling (see Section 5.3). Therefore, future work will focus on developing a comprehensive momentum balance correction model for the one-fluid formulation in a sharp VOF framework to include the correct treatment of the pressure jump due to viscous stresses and reduce the solution uncertainty.
@@
Additionally, this work has shown that the required resolution to achieve convergence can be very demanding in terms of cells per diameter of the bubble or droplet. This can be a result of having a very thin thermal boundary layer due to the properties of the fluid or it can be related to the resolution of surface tension. For practical simulations of multiple bubbles or droplets, realistic resolutions are much lower and require the development and integration of sub-models. Note that turbulence scales may be well-resolved to be considered a Direct Numerical Simulation framework, but under-resolution may still be an issue when it comes to interfacial phenomena. Thus, future work will consider heat and mass transfer sub-models for $\dot{m}^{\prime}$ and sub-models to accelerate, e.g., the bubble growth and necking formation in film boiling. Likewise, the momentum balance corrections might need some enhancement to converge faster to the expected pressure jump. Lastly, the proposed methodology must be extended to handle multiple objects, which includes the coupling with coalescence and breakup models.
@@ Appendix B Interface Embedding in the Discretization of the Energy Transport
A two-dimensional illustration of some possible interface intersections with the numerical stencils for the energy equation, Eq. (16), is given in Figure 1(a). The extension to three dimensions is straightforward. Rather than calculating the exact distance between a node and the intersection of the stencil with the interface, $\theta_{y}^{G}$, $\theta_{y}^{L}$ and $\theta_{x}^{L}$ are estimated from staggered volume fractions obtained geometrically by evaluating the enclosed volumes of two adjacent cells overlapping with the staggered cell [10, 16]. For the gas cell $(l,m)$ in Figure 1(a), the distance to the interface is given by $\theta_{y}^{G}\approx(1-C_{l,m-\frac{1}{2}})\Delta y$ and, for the liquid cell $(i,j)$, the distances are given by $\theta_{y}^{L}\approx C_{i,j+\frac{1}{2}}\Delta y$ and $\theta_{x}^{L}\approx C_{i-\frac{1}{2},j}\Delta x$. The steps to approximate $\theta_{y}^{G}$, $\theta_{y}^{L}$ and $\theta_{x}^{L}$ are visualized in Figure 1(b).
@@ Appendix C Time Step Calculation
$\Delta t_{\mathbf{u}}=\frac{2}{\tau_{\mathbf{u}}+\tau_{\mu}+\sqrt{(\tau_{\mathbf{u}}+\tau_{\mu})^{2}++4\tau_{\mathbf{g}}^{2}+4\tau_{\sigma}^{2}}}$ (49)
@@
$\begin{split}&\tau_{\mathbf{u}}=\frac{|u|_{\text{max}}}{\Delta x}+\frac{|v|_{\text{max}}}{\Delta y}+\frac{|w|_{\text{max}}}{\Delta z}\\ &\tau_{\mu}=2\bigg{(}\frac{1}{\Delta x^{2}}+\frac{1}{\Delta y^{2}}+\frac{1}{\Delta z^{2}}\bigg{)}\bigg{(}\frac{\mu}{\rho}\bigg{)}_{\text{max}}\\ &\tau_{\mathbf{g}}=\sqrt{\frac{||\mathbf{g}||}{\Delta z}}\\ &\tau_{\sigma}=\sqrt{\frac{(\sigma|\kappa|)_{\text{max}}}{\text{min}(\rho_{G},\rho_{L})\text{min}(\Delta x^{2},\Delta y^{2},\Delta z^{2})}}\end{split}$ (50)
@@
Additional time-step restrictions are given by the interface displacement due to phase change, $\Delta t_{\dot{m}^{\prime}}$, and the limitation of geometrical errors during advection of $C$ with split-advection solvers, $\Delta t_{C}$. These are given by limiting the maximum shift of the interface plane under phase change, $\Delta\text{d}_{\text{max}}$, as
$\Delta t_{\dot{m}^{\prime}}=\Delta\text{d}_{\text{max}}\frac{\rho_{L}}{\dot{m}^{\prime}}$ (52)
@@
$\Delta t=\text{min}\bigg{(}\theta_{1}\text{min}(\Delta t_{\mathbf{u}},\Delta t_{T}),\Delta t_{\dot{m}^{\prime}},\Delta t_{C}\bigg{)}$ (54)

### v2

@@ 6 Assessment of the Viscous Stress Jump with Evaporation
The increase in vapor volume $V$ normalized by the initial vapor volume $V_{0}$ and the average Nusselt number, Nu, on the bottom wall, defined as in Eq. (47) [17], are shown in Figure 21. Moreover, grid convergence is assessed with six mesh sizes (i.e., M1 to M6, with M6 corresponding to 1000 cells across $\lambda_{d}$), summarized in Table 6. The reference data shown here corresponds to a resolution across $\lambda_{d}$ of 128 cells for Guo et al. [61], 400 cells for Sun et al. [62], and 2048 cells for Boyd and Ling [17] (the latter using adaptive mesh refinement near the interface). Up to 0.2 s, the results agree well with the reference data and are fairly independent of mesh resolution. This corresponds to the initial growth of the vapor film before the instability accelerates and the bubble forms. For reference, Figure 22 shows the evolution of the mass-averaged vapor rise velocity and its center of mass obtained with M6 (i.e., the quantities are averaged over the vapor volume due to $\rho_{G}$ being constant). After that, substantial differences in $V/V_{0}$ and Nu are observed across different mesh resolutions and, more importantly, across different numerical frameworks.
@@
At first glance, one may claim that the solution from Guo et al. [61] is reproduced. However, upon closer inspection of the bubble shape, this is not the case. Instead, a shift in the time scale linked to the bubble formation, i.e., instability growth, is observed. The shape of the bubble is compared in Figure 23 against benchmark solutions at $t=0.42$ s [62, 17] and at $t=0.43$ s [61], and its growth is clearly slower. The plots also show the results with mesh M6 at $t=0.46$ s, which agree much better with the reference data. This delay between 0.03 s and 0.04 s is consistent with the shift in the $V/V_{0}$ and Nu profiles shown in Figure 21. When it comes to the numerical approach, the present work is close in methodology to Boyd and Ling [17] and considers the recent state-of-the-art in multiphase flow modeling. Some differences revolve around the conservative/non-conservative treatment of the momentum equation and the implementation of the volume dilatation due to phase change, but convergence of the numerical solution is observed with similar grid resolutions to M6. In fact, one could argue that earlier works are somewhat under-resolved [61, 62].
@@
Regardless of the exact onset, the numerical solution is very similar to previous literature. Thus, the differences must arise from the flow dynamics during the growth of the vapor film and the necking process leading to the bubble formation. Beyond the possibility that slight different initial conditions are affecting the time scales, the flow dynamics are affected by the specific numerical framework. These issues are not related to lack of convergence or to capturing the thermal layer as in Section 5.2. Figure 24 presents various snapshots of the solution at $t=0.15$ s, $t=0.25$ s, $t=0.35$ s and $t=0.45$ s obtained with M6, focusing on the contours of pressure, velocity magnitude, temperature and mass flux across the interface. It becomes immediately clear that a pressure spike occurs across the interface, which scales with $\dot{m}^{\prime\prime}$. This is a result of a momentum imbalance introduced by the discrete viscous term whereby “numerical regularization” fails. That is, a numerical pressure jump is introduced due to the calculation of $\nabla\mathbf{u}$ in the one-fluid framework, augmented by a high fluid viscosity. The thin region of the vapor film closer to the wall shows the largest momentum imbalance, while the error is much less significant along the bubble’s edge due to vapor cooling and the decrease in $\dot{m}^{\prime\prime}$. This regularization failure occurs independently of the type of mean used to average the viscosity, i.e., arithmetic vs. harmonic.
As depicted in Figure 24, the lateral pressure gradient that drives the vapor flow and bubble growth is a result of the varying film thickness and neck formation process. Because $\dot{m}^{\prime\prime}$ varies along the interface, the error introduced by the imbalance also does, modifying the pressure jump along the interface non-uniformly. Thus, the dynamics are easily affected, such as the spurious currents around the interface, and could explain the slower bubble growth with the current formulation. Note that despite the pressure spike across the interface scales with $\sim(\Delta x)^{-1}$, the solution seems to converge but to a wrong pressure field. This is visualized in Figure 25, showing the pressure extracted along $x=17$ mm at $t=0.45$ s.
@@
In this configuration, the impact of the momentum balance corrections from the Stefan flow shift and the addition of $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime\prime}}$ are minimal given that the viscous imbalance dominates. However, some differences are still observed that suggest that corrections in the treatment of the viscous term to improve consistency may help bring the solution closer to recent reference data [17] and reduce uncertainty. For example, Figure 26 shows a slight change in the shape of the bubble along its edge by considering or not the momentum balance corrections. The shape corresponds to the snapshot at $t=0.45$ s obtained with M6.
@@
A preliminary attempt to correct the viscous imbalance follows a phase-wise discretization of $\nabla\cdot\big{(}\mu\big{[}\nabla\mathbf{u}+(\nabla\mathbf{u})^{T}\big{]}\big{)}$. Here, phase-wise velocities are obtained by solving a Poisson equation in a narrow band around the interface to satisfy $\nabla\cdot\mathbf{u}_{l}=0$ and $\nabla\cdot\mathbf{u}_{g}=0$ [13]. Then, phase-wise values of velocity and viscosity are used at each face of the staggered velocity cell when discretizing the viscous term according to the predominant phase at the face. For that, the volume fraction in a cell centered around each face, $C_{\text{face}}$, is calculated geometrically. Liquid values are used if $C_{\text{face}}\geq 0.5$ and gas values are used if $C_{\text{face}}<0.5$. This discretization mitigates the momentum imbalance caused by the viscous term and improves the agreement with the instability timescale shown in previous work that may not suffer from the issue, as discussed in previous lines. This is shown in Figure 27, which plots again the evolution of the normalized vapor volume and the average Nu at the bottom wall obtained with the one-fluid and the phase-wise approaches using mesh M3.
Figure 27: Effects of the one-fluid vs. phase-wise discretization of the viscous term on the increase in vapor volume and the average Nusselt number at the wall of the two-dimensional film boiling case compared against benchmark solutions from Guo et al. [61], Sun et al. [62] and Boyd and Ling [17]. Mesh M3 is used. (a) vapor volume ratio $V/V_{0}$; (b) average Nusselt number.
@@ 7 Conclusions
A two-step correction has been proposed consisting of the addition of: (1) two body forces within the context of the CSF model [20], namely $\mathbf{f}_{NC}$ to correct or cancel the effects of an ill-defined convective term and $\mathbf{f}_{\dot{m}^{\prime\prime}}$ to recover the exact momentum jump due to phase change [12], and (2) a shift of the Stefan flow after advecting the interface to obtain an intermediate velocity field $\mathbf{u}^{*}$ before solving the momentum equation. The latter effectively addresses the momentum jump induced by the interface displacement in the non-conservative formulation [14], but imposes the assumption that the interface regression relative to the fluid is a quasi-steady process. Various tests have shown that these modifications successfully impose the correct momentum balance in viscous flows with low viscosity, recovering analytical solutions and reducing oscillations in the pressure field as the interface moves across grid cells. Moreover, the improvements are also evident in dynamic flows where the momentum balance corrections are seemingly negligible and have direct implications on the dynamic behavior of the flow. For example, bubble rise velocities may differ as shown in Section 5.2. However, the sharp discretization of the viscous term in highly viscous flows can induce pressure spikes across the interface and affect the momentum balance and dynamics of the flow, regardless of whether the conservative or non-conservative form of the momentum equation is used (see Section 6).
@@
It is in these scenarios that inconsistencies among different flow solvers are highlighted. The introduction of phase change is not trivial and requires a thorough examination of the impacts of a particular interface capturing scheme or discretization of the governing equations. Otherwise, different solvers will produce solutions similar in nature but with critical differences that may affect the particular analysis, e.g., the shift in the instability growth time scales seen in film boiling (see Section 6). Therefore, future work will focus on developing a comprehensive momentum balance correction model for the one-fluid formulation in a sharp VOF framework to include the correct treatment of the pressure jump due to viscous stresses and reduce the solution uncertainty.
@@
Additionally, this work has shown that the required resolution to achieve convergence can be very demanding in terms of cells per diameter of the bubble or droplet. This can be a result of having a very thin thermal boundary layer due to the properties of the fluid or it can be related to the resolution of surface tension. For practical simulations of multiple bubbles or droplets, realistic resolutions are much lower and require the development and integration of sub-models. Note that turbulence scales may be well-resolved to be considered a Direct Numerical Simulation framework, but under-resolution may still be an issue when it comes to interfacial phenomena. Thus, future work will consider heat and mass transfer sub-models for $\dot{m}^{\prime\prime}$ and sub-models to accelerate, e.g., the bubble growth and necking formation in film boiling. Likewise, the momentum balance corrections might need some enhancement to converge faster to the expected pressure jump. Lastly, the proposed methodology must be extended to handle multiple objects, which includes the coupling with coalescence and breakup models to prevent, e.g., numerical coalescence.
@@ Appendix B Interface Embedding in the Discretization of the Energy Transport
A two-dimensional illustration of some possible interface intersections with the numerical stencils for the energy equation, Eq. (16), is given in Figure 1(a). The extension to three dimensions is straightforward. Rather than calculating the exact distance between a node and the intersection of the stencil with the interface, $\theta_{y}^{G}$, $\theta_{y}^{L}$ and $\theta_{x}^{L}$ are estimated from staggered volume fractions obtained geometrically by evaluating the enclosed volumes of two adjacent cells overlapping with the staggered cell [48, 10, 16]. For the gas cell $(l,m)$ in Figure 1(a), the distance to the interface is given by $\theta_{y}^{G}\approx(1-C_{l,m-\frac{1}{2}})\Delta y$ and, for the liquid cell $(i,j)$, the distances are given by $\theta_{y}^{L}\approx C_{i,j+\frac{1}{2}}\Delta y$ and $\theta_{x}^{L}\approx C_{i-\frac{1}{2},j}\Delta x$. The steps to approximate $\theta_{y}^{G}$, $\theta_{y}^{L}$ and $\theta_{x}^{L}$ are visualized in Figure 1(b).
@@ Appendix C Time Step Calculation
$\Delta t_{\mathbf{u}}=\frac{2}{\delta t_{\mathbf{u}}+\delta t_{\mu}+\sqrt{(\delta t_{\mathbf{u}}+\delta t_{\mu})^{2}+4\delta t_{\mathbf{g}}^{2}+4\delta t_{\sigma}^{2}}}$ (49)
@@
$\begin{split}&\delta t_{\mathbf{u}}=\frac{|u|_{\text{max}}}{\Delta x}+\frac{|v|_{\text{max}}}{\Delta y}+\frac{|w|_{\text{max}}}{\Delta z}\\ &\delta t_{\mu}=2\bigg{(}\frac{1}{\Delta x^{2}}+\frac{1}{\Delta y^{2}}+\frac{1}{\Delta z^{2}}\bigg{)}\bigg{(}\frac{\mu}{\rho}\bigg{)}_{\text{max}}\\ &\delta t_{\mathbf{g}}=\sqrt{\frac{||\mathbf{g}||}{\Delta z}}\\ &\delta t_{\sigma}=\sqrt{\frac{(\sigma|\kappa|)_{\text{max}}}{\text{min}(\rho_{G},\rho_{L})\text{min}(\Delta x^{2},\Delta y^{2},\Delta z^{2})}}\end{split}$ (50)
@@
Additional time-step restrictions are given by the interface displacement due to phase change, $\Delta t_{\dot{m}^{\prime\prime}}$, and the limitation of geometrical errors during advection of $C$ with split-advection solvers, $\Delta t_{C}$. These are given by limiting the maximum shift of the interface plane under phase change, $\Delta\text{d}_{\text{max}}$, as
$\Delta t_{\dot{m}^{\prime\prime}}=\Delta\text{d}_{\text{max}}\frac{\rho_{L}}{\dot{m}^{\prime\prime}}$ (52)
@@
$\Delta t=\text{min}\bigg{(}\theta_{1}\text{min}(\Delta t_{\mathbf{u}},\Delta t_{T}),\Delta t_{\dot{m}^{\prime\prime}},\Delta t_{C}\bigg{)}$ (54)